I need help with the following problem:
Solve by the subsitution method.
9x + 2y = -37
-3x + y = 19

Answer by mananth(16946)

-3 x + 1 y = 19 .............1
1 y = 19 3 x ---------2
/ 1 =
y = 19 3 x
Substitute y in (2)
9 x + 2 ( 19 3 x ) =
9 x + 38 + 6 x = -37
9 x 6 x = -75
15 x = -75
/ 15
x= -5
Plug the value of x in (1)
9 x + 2 y = -37
9 * -5 + 2 y = -37
-45 + 2 y = -37
2 y = 8
/ 2
y= 4
m.ananth@hotmail.ca